# Zero-shot Class Unlearning via Layer-wise Relevance Analysis and Neuronal Path Perturbation

### v1

@@ IV-D Neuronal Path Perturbation
In neural networks for image classification, the initial layers typically extract low-level features such as edges and textures. The middle layers capture higher-level features like shapes and parts. In contrast, the FC layers converge the features into more abstract representations for classification so that we will perturb the FC layer $l_{i}$ for cutting off the neuronal classification path.
First, we extract the neurons’ weights that need to be perturbed. In Equation 11, we define the set of weights in layer $l_{i}$ that need to be perturbed as $W^{i}$. Formally, let $w^{i}$ represent the weights connecting layer $l_{i-1}$ to $l_{i}$. Each weight $w^{i}$ in $W^{i}$ corresponds with the position $p$ of neurons in $S^{i}$.
$W^{i}=\left\{w_{0}^{i},w_{1}^{i},...,w_{z-1}^{i}\right\},z\in\left[k,m\right]$ (11)
To facilitate the unlearning of a specific class from the neural network, we apply dropout to the neurons in layer $l_{i}$ responsible for classifying this class. As Equation 10 shows, $S^{i}$ denotes the set of neurons part of the neuronal classification path for the unlearning class in layer $l_{i}$. We implement dropout by setting the weights $w^{i}$, connecting the neurons in $S^{i}$ to the previous layer to 00 as Equation 12, effectively disrupting the classification path.
$w_{j}^{i}=0\quad\text{for all}\quad(w_{j}^{i}\in S^{i}\quad\text{and}\quad j\in\left[k,m\right])$ (12)
The neuronal path perturbation modifies the forward propagation with the dropout mask $M^{i}$ in $l_{i}$, where $M^{i}$ has entries of 0 for the weights corresponding to neurons in $S^{i}$ and $1$ otherwise, leading to the updated prediction function:
$F^{\prime}\left(x,\theta\right)=F\left(x,\theta\right)\cdot M^{i}$ (13)
By applying this targeted dropout, the network unlearns the specific class, as the classification path through layer $l_{i}$ is effectively cut off.
@@ IV-E Relationship Between Neuron Relevance and Unlearning
In deep learning models, neurons progressively extract and represent high-dimensional features of the input data through multi-layer structures and nonlinear activation functions. Neurons’ relevance to specific data samples reflects their role in processing these samples, indicating how well they capture the essential features of these samples. The relevance of neurons to data samples directly impacts the model’s classification performance on these samples.
Different neurons may have varying sensitivities to various features. Some neurons are highly sensitive to specific features that are key distinguishing characteristics of the data samples’ categories. Therefore, the activation states of these neurons significantly determine the model’s classification decisions. This implies that the model’s classification relies on neurons that can effectively extract and represent the input data’s features.
Identifying neurons with high relevance to the unlearning data becomes crucial in machine unlearning. By targeting and interfering with these specific neurons, we can effectively disrupt their ability to capture and represent the features of the unlearning data. This disruption can be achieved by altering the weights or applying noise to the outputs of these neurons. Consequently, the model’s ability to classify the unlearning data will be impaired, achieving the goal of unlearning.
This targeted interference should be designed to affect the neurons’ role in processing minimally and classifying the remaining data. Because the neurons with high relevance to the unlearning data are specifically identified and targeted, the model’s performance on the remaining data can be preserved. This selective approach ensures that the model maintains its accuracy and robustness for the data not subject to unlearning, thereby achieving effective unlearning without compromising the overall model performance.
V Expriment
V-A Datasets and Setups
In our experiments, we utilized various datasets and models. We will introduce our work separately for image classification tasks.
MNIST Dataset [24]. The MNIST dataset is a large database of handwritten digits commonly used for training various image processing systems. The dataset contains $60,000$ training images and $10,000$ testing images, each with a resolution of $28$x28282828 pixels.
CIFAR-10 Dataset [25]. The CIFAR-10 dataset consists of $60,000$ color images in $10$ different classes, with $6,000$ images per class. Each image has a resolution of $32$x32323232 pixels and is widely used for training and evaluating machine learning algorithms in image classification.
CIFAR-100 Dataset [25]. The CIFAR-100 dataset is similar to CIFAR-10 but contains $100$ classes, each with $600$ images, making a total of $60,000$ images. The images are also $32$x32323232 pixels in size, and this dataset is often used for more challenging tasks in image recognition, requiring algorithms to learn finer distinctions among categories.
mini-ImageNet Dataset [26]. mini-ImageNet is a massive database of images that are hand-annotated and categorized into $1,000$ different classes, ranging from objects to scenes to abstract concepts. It’s like an enormous visual dictionary for computers, particularly in image recognition and classification.
@@ V-A Datasets and Setups
Our experiments preprocessed the CIFAR-10, CIFAR-100, and mini-ImageNet datasets to adjust their resolutions, making them more aligned with real-world image scenarios. Table II provides each dataset’s image resolution and the corresponding classification model. Furthermore, to verify that our method can address Challenge $2$ by enabling the model owner to unlearn the unlearning class without accessing the training data, we used data from the test set of each dataset as samples for layer-wise relevance analysis.
@@ V-B1 Evaluation Metrics for Images unlearning
In our experiments on image unlearning, we utilized a range of intuitive and precise evaluation metrics to illustrate the effectiveness of our proposed approach. The metrics employed are outlined below:
Unlearning class accuracy ($A_{t}$). The accuracy of the model on the unlearning class, $A_{t}$, reflects the model’s performance during the forgetting process. A high $A_{t}$ suggests that the model struggles to remove the knowledge of the target class from its memory. Conversely, a low $A_{t}$ indicates successful unlearning to the target class.
Global accuracy ($A_{g}$). The model’s average accuracy on the remaining classes is denoted as $A_{g}$. $A_{g}$ can be obtained by calculating the model’s accuracy on the dataset after removing the data from the target unlearned class. A high $A_{g}$ reflects the improved utility of the unlearned model for the remaining tasks. Furthermore, a lower $A_{g}$ signifies a decrease in the model’s utility after unlearning.
Forgetting rate ($Fr$). The forgetting rate of the unlearned model is demonstrated as $Fr$. $Fr$ is calculated by the proportion of ”data not trained by the model” in the results of the membership inference attack. A higher $Fr$ indicates that more unlearning data has been forgotten.
Gigabyte ($G$). In this context, Gigabyte (GB) is employed to measure the computational resource consumption of GPU memory. A Gigabyte represents $10^{9}$ bytes of data. We denote Gigabyte as $G$. Higher $G$ signifies higher resource consumption.
Time. The time spent by the unlearning process is demonstrated as $T$. $T$ records the time taken from the model owner’s initiation of unlearning until completion. The method requires less $T$ for machine unlearning, indicating its greater efficiency.
@@ V-B2 Comparing methods list
  1. 1.
    Retrain from Scratch involves completely retraining the model from its initial state after removing specific data points. This method ensures that the influence of the removed data is entirely eradicated from the model.
  2. 2.
    Random Labels [27] using random labels for unlearning involves randomizing the data labels of the target class, making it difficult for the model to recognize and learn the features of these data, thereby achieving effective unlearning of that class.
  3. 3.
    Amnesiac Unlearning [28] involves tracking the parameter updates from each batch during training and, upon receiving a data removal request, selectively undoing the parameter updates from the batches containing the sensitive data, thereby efficiently removing specific learned data from the neural network model.
  4. 4.
    Boundary Shrink [4] achieves machine unlearning by identifying and targeting the decision boundary of the class to be forgotten. It reassigns labels of unlearning samples to nearby incorrect classes, shrinking the decision space for the unlearning class.
  5. 5.
    Boundary Expanding [4] facilitates rapid unlearning by introducing a shadow class to disperse the activations of unlearning samples within an expanded decision space. After finetuning, the shadow class is pruned, removing the influence of unlearning data.
  6. 6.
    Neuronal Path Noise refers to introducing Gaussian or Laplacian noise to neurons identified as relevant to the unlearning dataset using the LRP method. This added noise aims to disrupt the model’s ability to classify the unlearning dataset accurately.
@@ V-C1 Changes in quantitative metrics
Information about specific classes is relatively easy to obtain in image classification tasks. We employ the proposed unlearning method to unlearn a specific class in the MNIST, CIFAR-10, CIFAR-100, and mini-ImageNet datasets.
We choose retraining as a baseline for unlearning. The performance of the retrained model in the unlearning class provides a baseline for assessing the effectiveness of unlearning methods. Then, We will apply the $\epsilon$-rule to LRP to analyze neuron relevance to verify the unlearning method’s effectiveness. $\epsilon$-rule involves adding a small term $\epsilon$ to the denominator in the relevance propagation formula to avoid zero division and control the redistribution of relevance scores.
@@
TABLE III: Class Unlearning Performance on the MNIST dataset
@@
Retrain from scratch $0.00$ $0.97$ $1.00$ $0.74$ $71.37s$
Random labels $0.00$ $0.97$ $1.00$ $0.70$ $15.64s$
@@
Boundary expanding $0.13$ $0.95$ $0.91$ $0.65$ $69.21s$
Boundary shrink $0.05$ $0.58$ $1.00$ $0.71$ $18.55s$
Neuronal path noise (GN) $0.35$ $0.68$ $0.61$ $0.64$ $0.77s$
Neuronal path noise (LN) $0.21$ $0.53$ $0.78$ $0.64$ $0.83s$
Neuronal path perturbation $0.00$ $0.97$ $1.00$ $0.64$ $0.68s$
In Table III, after applying our unlearning method, the model’s accuracy $A_{t}$ on class $1$ dropped to $0.00$, matching the $A_{t}$ observed after retraining. This demonstrates that the model has completely lost its ability to classify data from the unlearning class. Disrupting the neurons associated with the unlearning class makes it possible to cut off the neural network’s classification path.
While cutting off the model’s classification path for the unlearning class, we preserve the model’s ability to classify the remaining classes. For example, after retraining from scratch, the model achieves an overall accuracy $A_{g}$ of $0.97$ on the remaining classes. By applying our machine unlearning method, the model maintains an $A_{g}$ of $0.97$ on these classes, demonstrating strong performance. This shows that our method effectively unlearns specific class information without additional training and preserves the model’s classification capability.
Our method also excels in unlearning verification through membership inference attacks. After retraining, the new model achieves a perfect forgetting rate of $1.00$ for the unlearning class, indicating that it is entirely unaffected by the unlearning data. Our approach also achieves a forgetting rate of $1.00$, demonstrating that severing the classification path in the neural network effectively accomplishes the class unlearning task.
Also, our method performs well in terms of computational resource consumption. When using retraining to achieve unlearning, the GPU memory consumption $G$ was $0.74$, which took $71.37$ seconds. In contrast, we completed class unlearning in just $0.68$ seconds using our unlearning method. Additionally, GPU resource consumption was only $0.64$, and we used just $36$ images from class $1$ for unlearning. These results show that our unlearning method reduces resource consumption without additional training.
@@
Retrain from scratch $0.00$ $0.96$ $1.00$ $6.89$ $484.17s$
Random labels $0.00$ $0.90$ $0.99$ $6.89$ $107.99s$
@@
Neuronal path noise (GN) $0.95$ $0.79$ $0.00$ $2.82$ $14.17s$
Neuronal path noise (LN) $0.99$ $0.71$ $0.00$ $2.82$ $15.51s$
Neuronal path perturbation $0.03$ $0.97$ $0.99$ $2.82$ $14.44s$
The results in Table IV demonstrate that Neuronal Path Perturbation is the most effective method for class unlearning on the CIFAR-10 dataset, achieving an excellent balance between unlearning performance and model utility. With an accuracy of $A_{t}=0.03$ on the target class, it ensures nearly complete unlearning, comparable to retraining from scratch ($A_{t}=0.00$) and random labels ($A_{t}=0.00$). Its forgetting rate ($Fr=0.99$) is almost as high as the ideal standard of $Fr=1.00$, indicating a solid unlearning effect.
At the same time, Neuronal Path Perturbation maintains high accuracy on the remaining classes ($A_{g}=0.97$), outperforming most other methods and closely matching the retrained model’s $A_{g}=0.96$. This reflects its ability to retain model utility while effectively unlearning the target class.
Moreover, Neuronal Path Perturbation is highly efficient in terms of computational resources. It consumes only $2.82$ GB of GPU memory, significantly less than retraining ($6.89$ GB), and takes just $14.44$ seconds to complete the unlearning process, a fraction of the $484.17$ seconds required for retraining. This makes it a highly practical and efficient choice for unlearning tasks.
@@
Amnesiac $0.03$ $0.70$ $0.29$ $1.75$ $18.45s$
Boundary expanding $0.77$ $0.81$ $0.00$ $4.07$ $83.36s$
@@
Neuronal path noise (LN) $0.00$ $0.29$ $0.00$ $2.83$ $0.37s$
Neuronal path perturbation $0.01$ $0.82$ $1.00$ $2.77$ $0.53s$
The results for class unlearning on the CIFAR-100 dataset highlight that our method, Neuronal Path Perturbation, offers the best balance between achieving effective unlearning and preserving model utility. With an accuracy of $A_{t}=0.01$ on the target class, this method achieves near-complete unlearning, comparable to retraining from scratch and random labels (both $A_{t}=0.00$). Its forgetting rate ($Fr=1.00$) is identical to these methods, eliminating the learned information about the target class.
Despite this solid unlearning performance, Neuronal Path Perturbation preserves a high level of accuracy on the remaining classes ($A_{g}=0.82$), matching that of the retrained model. This shows that the method effectively retains the model’s utility across the non-target classes, making it highly efficient in maintaining overall performance.
Regarding resource efficiency, Neuronal Path Perturbation consumes only $2.77$ GB of GPU memory, much lower than retraining ($6.89$ GB). It requires just $0.53$ seconds to complete the unlearning process—an impressive speed compared to the $607.89$ seconds for retraining. These results make it a highly efficient and practical solution for class unlearning tasks on CIFAR-100.
@@
Random labels $0.17$ $0.14$ $0.74$ $2.47$ $582.32s$
Amnesiac $0.53$ $0.80$ $0.44$ $7.61$ $20.09s$
Boundary expanding $0.03$ $0.64$ $1.00$ $7.68$ $697.89s$
Boundary shrink $0.35$ $0.67$ $0.76$ $9.13$ $4.03s$
@@
Neuronal path perturbation $0.00$ $0.83$ $1.00$ $0.93$ $1.35s$
For a VGG-$16$ model pre-trained on the ImageNet dataset, we employed machine unlearning using $33$ images related to the unlearning class. The mini-ImageNet dataset also has $1,000$ classes. While each class has fewer images than ImageNet, it’s still a good way to test if unlearning methods work for real-world multi-class tasks. Our experimental purpose is to make the pre-trained VGG model forget the knowledge about class 00.
As shown in Table VI, after applying our unlearning method, the model’s $A_{t}$ dropped to $0.00$, while $A_{g}$ remained at $0.83$. Our method’s unlearning results are consistent with those obtained from validating the retrained model. These results indicate that our method enabled the model to unlearn information specific to the unlearning class.
Similarly, employing a membership inference attack based on SVM for unlearning verification on VGG yielded favorable results. After retraining the model, the $Fr$ is $1.00$. Moreover, the $Fr$ of the unlearned model obtained using our method is also $1.00$, indicating that the model has wholly unlearned the knowledge of the unlearning class data. The results demonstrate that perturbing the neuronal classification path is sufficient to make the model lose its classification ability for a specific class.
When retraining the model, we set the training batch size to $32$ to achieve better results. The retraining process consumed approximately $6.51$G GPU memory. In contrast, our method only involves model inference, consuming just $0.93$G GPU memory, significantly less than the computational resources required for retraining. Moreover, from the time cost perspective, our method only requires $0.68$ seconds to complete unlearning, whereas retraining takes $2865.87$ seconds. This demonstrates that our method is highly efficient.

### v2

@@ IV-D Neuronal Path Perturbation
In neural networks for image classification, the initial layers typically extract low-level features such as edges and textures. Middle layers capture higher-level features like shapes and parts. In contrast, the FC layers transform these features into more abstract representations for classification. Therefore, we perturb the FC layer $l$ to disrupt the neuronal classification path.
First, we identify the weights connected to the neurons that need to be perturbed. These are the weights connecting the preceding layer $l-1$ to the selected neurons in layer $l$. As identified by the relevance analysis in the previous section, the set of strongly relevant neuron positions in layer $l$ for the unlearned class is $S_{l}$. In Equation 10, we represent the set containing all incoming weights connected to the neurons identified in $S_{l}$ as $W_{l}$.
$W_{l}=\left\{w_{0},w_{1},...,w_{z-1}\right\},z\in\left[k,m\right]$ (10)
To facilitate unlearning of a specific class, we apply targeted dropout by modifying the weights connected to the neurons in layer $l$ responsible for classifying this class. As Equation 9 shows, $S_{l}$ is the list of strongly relevant neuron positions part of the neuronal classification path for the unlearning class in layer $l$. We implement this by setting the incoming weights to the neurons whose positions are in $S_{l}$ to 00. Specifically, for each neuron position $p\in S_{l}$, all weights connecting from layer $l-1$ to the neuron at position $p$ in layer $l$ are set to 00. This effectively disrupts the classification path.
This neuronal path perturbation modifies the forward propagation. We implement this by setting the weights corresponding to neurons in $S_{l}$ to 00. This creates modified network parameters $\theta^{\prime}$, where the weights in layer $l$ connected to neurons in $S_{l}$ are zeroed out. The updated prediction function is then computed using these modified parameters, denoted as $F^{\prime}\left(x\right)=F\left(x,\theta^{\prime}\right)$.
By applying this targeted weight masking (or dropout), the network unlearns the specific class, as the classification path through layer $l$ is effectively severed.
@@ IV-E Relationship Between Neuron Relevance and Unlearning
We demonstrate that by disabling key neurons highly relevant to a specific class in a neural network, we can effectively make the network unlearn that class. To identify these neurons, we need relevance scores to analyze the weights of neurons within the neural network. This leads to the following claim:
Claim 1
Setting the weights $w_{p}$ of neurons $p$ with high relevance scores $R_{p}(x_{u})$ for the unlearning class $C_{i}$ to zero effectively achieves unlearning of class $C_{i}$.
This operation results in the model’s prediction function $F(x;\theta)$ losing accuracy for classifying data from class $C_{i}$, thus realizing effective unlearning. Specifically, for an input sample $x_{u}$ belonging to class $C_{i}$, the weights are adjusted as follows:
$w_{p}=\begin{cases}0&\text{if }R_{p}(x_{u})\gg R_{q}(x_{u})\text{ for all }q\notin\{p\}_{\text{relevant},u}\\ w_{p}&\text{otherwise}\end{cases}$ (11)
where $\{p\}_{\text{relevant},u}$ is the set of neurons with high relevance scores for the input sample $x_{u}$.
This adjustment modifies the model’s prediction function as shown in Equation 12, where $f(\cdot)$ is the activation function applied at the final layer, and $a$ represents the activation values.
$F(x;\theta^{\prime})$ $=f\left(\sum_{\begin{subarray}{c}p\in\{p\}_{\text{relevant},u}\\ R_{p}(x_{u})\gg R_{q}(x_{u})\end{subarray}}0\cdot a_{p}+\sum_{q\notin\{p\}_{\text{relevant},u}}w_{q}a_{q}\right)$ (12)
$=f\left(\sum_{q\notin\{p\}_{\text{relevant},u}}w_{q}a_{q}\right)$
To prove this, consider the effect of changes in the activations of the highly relevant neurons for a given input $x$, denoted as the set $\{p\}_{\text{relevant},x}$, on the model output $F(x;\theta)$. The output $F(x;\theta)$ can be expressed as a function of the activations of both relevant and irrelevant neurons, as stated in Claim 2.
Claim 2
The output $F(x;\theta)$ is a weighted sum of the activations $a_{p}$ of both relevant and irrelevant neurons.
The output is then passed through an activation function $f(\cdot)$ to introduce non-linearity. Specifically, the output can be expressed as Equation 13, where $w_{p}$ and $a_{p}$ represent the weight and activation associated with neuron $p$.
$F(x;\theta)=f\left(\sum_{p\in\{p\}_{\text{relevant},x}}w_{p}a_{p}+\sum_{p\in\{p\}_{\text{irrelevant},x}}w_{p}a_{p}\right)$ (13)
where $\{p\}_{\text{irrelevant},x}$ represents the set of neurons with low relevance scores for the input $x$.
Let’s now analyze the effect of perturbing the activations of the relevant neurons for a given input $x$. We first define the change in the model output as follows:
Claim 3
The change in the model output $\Delta F(x;\theta)$ is caused by the change in the activations $\Delta a_{p}$ of the neurons, and the magnitude of this change is determined by the sensitivity of the neuron, $\frac{\partial F(x;\theta)}{\partial a_{p}}$.
Based on Claim 2, if we introduce a small change $\Delta a_{p}$ to each activation $a_{p}$ for $p\in\{p\}_{\text{relevant},x}$, the resulting change in the output $\Delta F(x;\theta)$ can be approximated using a first-order Taylor expansion around the initial activations:
$\Delta F(x;\theta)\approx$ $\sum_{p\in\{p\}_{\text{relevant, x}}}\frac{\partial F(x;\theta)}{\partial a_{p}}\Delta a_{p}$
$+\frac{1}{2}\sum_{p,q\in\{p\}_{\text{relevant, x}}}\frac{\partial^{2}F(x;\theta)}{\partial a_{p}\partial a_{q}}\Delta a_{p}\Delta a_{q}+\cdots$ (14)
For small perturbations $\Delta a_{p}$, higher-order terms can be ignored, leaving the linear approximation:
$\Delta F(x;\theta)\approx\sum_{p\in\{p\}_{\text{relevant},x}}\frac{\partial F(x;\theta)}{\partial a_{p}}\Delta a_{p}$ (15)
where $\frac{\partial F(x;\theta)}{\partial a_{p}}$ is the sensitivity of the output to the activation $a_{p}$ of neuron $p$ for the input $x$. Since neurons in $\{p\}_{\text{relevant},x}$ have high relevance scores $R_{p}(x)$, they typically exhibit large values for $\frac{\partial F(x;\theta)}{\partial a_{p}}$, meaning that even small changes $\Delta a_{p}$ can lead to significant changes in $F(x;\theta)$.
Thus, we have:
$|\Delta F(x;\theta)|=\left|\sum_{p\in\{p\}_{\text{relevant},x}}\frac{\partial F(x;\theta)}{\partial a_{p}}\Delta a_{p}\right|\gg 0$ (16)
Equation 16 demonstrates that the model’s output is highly sensitive to changes in the activations of neurons with high relevance scores $R_{p}(x)$ for the input $x$. Consequently, these relevant neurons play a crucial role in determining the accuracy of the model’s classification performance on the sample $x$.
Claim 4
Specific neurons store key information about a class, leading to their high relevance to the data of that class. High-relevance neurons for class $C_{i}$, denoted as $\{p\}_{C_{i}}$, can be distinguished from low-relevance ones based on their significant relevance scores $R_{p}(x)$, where $R_{p}(x)\gg R_{q}(x)$ for $p\in\{p\}_{C_{i}}$ and $q\notin\{p\}_{C_{i}}$ for data $x$ belonging to class $C_{i}$.
Different neurons exhibit varying sensitivities to features, with some neurons having higher relevance scores $R_{p}(x)$ for key characteristics of specific categories. These neurons, $p\in\{p\}_{C_{i}}$, are crucial for classification, and we may observe some overlap between categories:
$\{p\}_{C_{i}}\cap\{p\}_{C_{j}}\neq\emptyset,\quad\forall C_{i},C_{j}\in\mathcal{C},\,i\neq j$ (17)
Therefore, identifying neurons highly relevant to the unlearning data $x_{u}$ is essential for machine unlearning. For a sample $x_{u}$ belonging to the unlearning class, let $\{p\}_{\text{relevant},u}$ represent the subset of neurons with high relevance scores $R_{p}(x_{u})$. These neurons are crucial for capturing and representing the key features of $x_{u}$, as expressed by:
$R_{p}(x_{u})\gg R_{q}(x_{u})\quad\text{for }p\in\{p\}_{\text{relevant},u},\;q\notin\{p\}_{\text{relevant},u}$ (18)
By identifying the neurons in $\{p\}_{\text{relevant},u}$, we locate those primarily responsible for encoding information about the unlearning data.
Based on Claim 2 to 4, a subset of neurons with high relevance plays a key role in the classification of a specific class. Therefore, to achieve Claim 1, we set the weights $w_{p}$ of these highly relevant neurons for the unlearning sample $x_{u}$ to zero, effectively removing their contributions to the model’s output for that sample. Substituting $w_{p}=0$ for $p\in\{p\}_{\text{relevant},u}$ in the output function:
$F(x;\theta^{\prime})$ $=f\left(\sum_{p\in\{p\}_{\text{relevant},u}}0\cdot a_{p}+\sum_{p\in\{p\}_{\text{irrelevant},u}}w_{p}a_{p}\right)$
$=f\left(\sum_{p\in\{p\}_{\text{irrelevant},u}}w_{p}a_{p}\right)$ (19)
where $\{p\}_{\text{irrelevant},u}$ represents the set of neurons with low relevance scores for the input sample $x_{u}$.
V Experiment
V-A Datasets and Models
@@
We aaplied MNIST, CIFAR-10, CIFAR-100, and mini-ImageNet datasets for evaluating the performance of our method. Our experiments preprocessed the CIFAR-10, CIFAR-100, and mini-ImageNet datasets to adjust their resolutions, making them more aligned with real-world image scenarios. Table II provides each dataset’s image resolution and the corresponding classification model. Furthermore, to ensure that our method satisfies the definition of “Zero-shot Unlearning,” which requires the model owner to unlearn a class without accessing the training data, we extract data from the test set of each dataset as “unseen” samples for layer-wise relevance analysis.
@@ V-B1 Evaluation Metrics for Images unlearning
  • Unlearning class accuracy ($A_{t}$). The accuracy of the model on the unlearning class. A low $A_{t}$ indicates successful unlearning to the target class. This metric directly quantifies the model’s reduced ability to correctly classify the data points intended to be forgotten.
  • Global accuracy ($A_{g}$). The model’s average accuracy on the remaining classes is denoted as $A_{g}$. A high $A_{g}$ reflects the improved utility of the unlearned model for the remaining tasks. Maintaining a high $A_{g}$ ensures that the unlearning process does not significantly degrade the model’s performance.
  • Forgetting rate ($Fr$). The forgetting rate ($Fr$) of the unlearned model is a metric derived from a membership inference attack (MIA). This attack involves training an SVM attack model using the target model’s posterior probabilities from its training data to learn how to identify unlearned class members. Subsequently, this trained SVM is used to classify the posterior probabilities produced by the unlearned model when applied to the unlearned class data. Based on this classification result, $Fr$ is calculated as the proportion of unlearned data points that this attack fails to identify as members. A higher $Fr$ value indicates more successful forgetting.
  • Gigabyte ($G$). In this context, Gigabyte is employed to measure the unlearning computational resource consumption of GPU memory. We denote Gigabyte as $G$. A higher $G$ signifies a higher resource consumption. Specifically, $G$ quantifies the peak GPU memory required by the unlearning algorithm during its execution.
  • Time. The time spent by the unlearning process is demonstrated as $T$. A less $T$ indicates greater efficiency. This measures the total wall-clock time required for the unlearning algorithm to complete.
@@ V-B2 Comparing methods list
  • Retrain from Scratch: Retrains the model from scratch after removing specific data, completely erasing the influence of the removed data.
  • Random Labels [23]: Randomizes the target class labels, making it harder for the model to recognize and learn from those data, achieving unlearning.
  • Amnesiac Unlearning [24]: Tracks parameter updates during training and selectively undoes updates from batches containing sensitive data to remove its influence.
  • Boundary Shrink [5]: Targets the decision boundary by reassigning unlearning samples to incorrect classes, shrinking the decision space of the target class.
  • Boundary Expanding [5]: Expands the decision space by introducing a shadow class to disperse activations, then prunes it after finetuning to remove unlearning data’s influence.
  • Neuronal Path Noise: Introduces Gaussian (GN) or Laplacian (LN) noise to neurons identified as relevant to the unlearning dataset, disrupting the model’s ability to classify those data accurately. When applying Gaussian noise, the parameters were set to a mean of 0.0 and a standard deviation of 1.0; for Laplace noise, the parameters were set to a location of 0.0 and a scale parameter of 1.0.
@@ V-C Performance Evaluation
In image classification tasks, we apply the proposed unlearning method to remove a specific class from the MNIST, CIFAR-10, CIFAR-100, and mini-ImageNet datasets. Figures 2 to 5 show the attribution results of LRP for unlearning and non-unlearning data before and after machine unlearning.
@@ V-C1 Changes in quantitative metrics
We use retraining as the baseline for unlearning. The retrained model’s performance on the target class serves as a reference for evaluating unlearning effectiveness. To assess the method, we apply the $\epsilon$-rule in LRP to analyze neuron relevance. In Table III to VI, for each column, the bold and underlined values represent the best and second-best results respectively. Our method delivers the most effective and well-balanced results.
@@
TABLE III: Class Unlearning Performance on the MNIST dataset.
@@
Retrain from scratch 0.00 $\underline{0.97}$ 1.00 $0.74$ $71.37s$
Random labels 0.00 $\underline{0.97}$ 1.00 $0.70$ $15.64s$
@@
Boundary expanding $0.13$ $0.95$ $\underline{0.91}$ $0.65$ $69.21s$
Boundary shrink $\underline{0.05}$ $0.58$ 1.00 $0.71$ $18.55s$
Neuronal path noise (GN) $0.35$ $0.68$ $0.61$ $\underline{0.64}$ $\underline{0.77s}$
Neuronal path noise (LN) $0.21$ $0.53$ $0.78$ $\underline{0.64}$ $0.83s$
Neuronal path perturbation 0.00 $\underline{0.97}$ 1.00 $\underline{0.64}$ 0.68s
In Table III, after applying our unlearning method, the model’s accuracy $A_{t}$ on class 1 dropped to $0.00$, matching the $A_{t}$ after retraining. This shows the model completely loses the ability to classify data from the unlearning class, cutting off the classification path by disrupting relevant neurons.
While removing the unlearning class path, the model’s ability to classify other classes is preserved. After retraining, the model achieves an overall accuracy $A_{g}$ of $0.97$ on the remaining classes. With our unlearning method, $A_{g}$ remains $0.97$.
Our method also performs well in unlearning verification through membership inference. After retraining, the new model achieves a forgetting rate of $1.00$ for the unlearning class, indicating full unlearning. Our approach achieves the same forgetting rate, confirming that severing the classification path effectively unlearns the class.
Moreover, our method is computationally efficient. Retraining requires $71.37$ seconds and $0.74$ GPU memory, while our unlearning method completes the task in $0.68$ seconds with only $0.64$ GPU memory. These results show that our method reduces resource consumption without additional training.
@@
Retrain from scratch 0.00 $\underline{0.96}$ 1.00 $6.89$ $484.17s$
Random labels 0.00 $0.90$ $\underline{0.99}$ $6.89$ $107.99s$
@@
Neuronal path noise (GN) $0.95$ $0.79$ $0.00$ $\underline{2.82}$ $\underline{14.17s}$
Neuronal path noise (LN) $0.99$ $0.71$ $0.00$ $\underline{2.82}$ $15.51s$
Neuronal path perturbation $\underline{0.03}$ 0.97 $\underline{0.99}$ $\underline{2.82}$ $14.44s$
The results in Table IV show that Neuronal Path Perturbation is the most effective unlearning method for the CIFAR-10 dataset, achieving near-complete unlearning with $A_{t}=0.03$, similar to retraining ($A_{t}=0.00$). Its forgetting rate ($Fr=0.99$) is close to the ideal $Fr=1.00$, indicating strong unlearning.
At the same time, it maintains high accuracy on the remaining classes ($A_{g}=0.97$), outperforming most methods and matching retraining ($A_{g}=0.96$). This demonstrates its ability to preserve model utility while effectively unlearning the target class.
Additionally, Neuronal Path Perturbation is highly efficient, using only $2.82$ GB of GPU memory and taking just $14.44$ seconds, compared to retraining’s $6.89$ GB and $484.17$ seconds, making it a practical and efficient choice for unlearning.
@@
Amnesiac $0.03$ $0.70$ $\underline{0.29}$ 1.75 $18.45s$
Boundary expanding $0.77$ $\underline{0.81}$ $0.00$ $4.07$ $83.36s$
@@
Neuronal path noise (LN) 0.00 $0.29$ $0.00$ $2.83$ $\underline{0.37s}$
Neuronal path perturbation $\underline{0.01}$ 0.82 1.00 $\underline{2.77}$ $0.53s$
The CIFAR-100 class unlearning results show that Neuronal Path Perturbation strikes the best balance between effective unlearning and model utility. With $A_{t}=0.01$ on the target class, it achieves near-complete unlearning, matching retraining and random labels ($A_{t}=0.00$). Its forgetting rate ($Fr=1.00$) fully erases target class information.
At the same time, it maintains high accuracy on remaining classes ($A_{g}=0.82$), matching the retrained model. This demonstrates its ability to preserve model utility while unlearning the target class.
In terms of resource efficiency, Neuronal Path Perturbation uses only $2.77$ GB of GPU memory, far less than retraining ($6.89$ GB), and completes the process in $0.53$ seconds, compared to retraining’s $607.89$ seconds. This makes it a highly efficient and practical solution for CIFAR-100 class unlearning.
@@
Random labels $0.17$ $0.14$ $0.74$ $\underline{2.47}$ $582.32s$
Amnesiac $0.53$ $\underline{0.80}$ $0.44$ $7.61$ $20.09s$
Boundary expanding $\underline{0.03}$ $0.64$ 1.00 $7.68$ $697.89s$
Boundary shrink $0.35$ $0.67$ $\underline{0.76}$ $9.13$ $4.03s$
@@
Neuronal path perturbation 0.00 0.83 1.00 0.93 $\underline{1.35s}$
For the VGG-16 model pre-trained on ImageNet, we applied machine unlearning using $33$ images from the unlearning class on the mini-ImageNet dataset, which has $1,000$ classes. The goal was to make the model forget class 00.
As shown in Table VI, after applying our method, the model’s $A_{t}$ dropped to $0.00$, while $A_{g}$ remained at $0.83$, matching the results of retraining. This confirms that our method effectively unlearned class-specific knowledge.
Verification with a membership inference attack using SVM also showed a $Fr$ of $1.00$ after both retraining and our unlearning method, indicating complete unlearning of the target class.
Retraining with a batch size of $32$ used $6.51$ GB of GPU memory, while our method only required $0.93$ GB. Additionally, our method completed unlearning in $0.68$ seconds, compared to $2865.87$ seconds for retraining, highlighting its efficiency.